# Results of Fractional Rough Burgers equation in $H^{s}$ space and its application

### v1

@@ 3.2 Well-posedness
be satisfied for some $\sigma,s>0$, $p>1$. Note that $\partial_{t}v=\Lambda^{\gamma}v+\frac{1}{2}\partial_{x}(v^{2})+\partial_{x}(vX)+\frac{1}{2}\partial_{x}(X^{2})$. By Lemma 2.1 and Proposition 3.2, we have $\partial_{t}v\in L_{T}^{2}H^{-\sigma}$ for any $\sigma\geq 1$. The second condition of (3.13) is a consequence of (3.4). Then we prove $v$ is the solution of (3.3) which prove our claim.
To obtain the global solution, we employ induction argument. Fix $T^{*}\in[0,\infty)$, let $T^{(0)}\leq T^{*}$ small enough satisfied right hand side of (3.15) less than $\|u_{0}\|_{L^{2}}$($\tilde{T}$ can be only dependent on $\|X\|_{C_{T^{*}}W^{\alpha}}$). Consider equation which is begin at $v(\frac{T^{(0)}}{2})$. Fix $v^{(1)}=v(\frac{T^{(0)}}{2}+t)$, $X^{(1)}=X(\frac{T^{(0)}}{2}+t)$ then $v^{(1)}(t)$ satisfied the equation
$\partial_{t}v^{(1)}(t)-\Lambda^{\gamma}v^{(1)}(t)=-\partial_{x}(\frac{1}{2}(v^{(1)}+X^{(1)})^{2}),\quad v^{(1)}(0)=v(\frac{T^{(0)}}{2}).$ (3.14)
then we can get a solution $v^{(1)}$ of equation on $\frac{T^{(0)}}{2},\frac{3T^{(0)}}{2}]$ since $\|v(\frac{T}{2})\|_{L^{2}}\leq\|u_{0}\|_{L^{2}}<c$ and
$\|X^{(1)}\|_{C_{T}(W^{\frac{1}{4}-\kappa})}=\|X\|_{C_{[\frac{T^{(0)}}{2},\frac{T^{(0)}}{2}+T]}W^{\frac{1}{4}-\kappa}}\lesssim\|X\|^{2}_{C_{T^{*}}W^{\frac{1}{4}-\kappa}}.$
$v^{(1)}$ on $\frac{T^{(0)}}{2},\frac{3T^{(0)}}{2}]$ is also a solution of (3.3). Then define $v^{new}:=v^{(0)}$ on $0,\frac{T^{(0)}}{2}]$ and $v^{new}:=v^{(1)}$ on $\frac{T^{(0)}}{2},\frac{3T^{(0)}}{2}]$. Then $v^{new}$ is a solution in $C([0,\frac{3T^{(0)}}{2}];L^{2}(\mathbb{T}))\cap L^{2}([0,\frac{3T^{(0)}}{2}];H^{\frac{\gamma}{2}}(\mathbb{T}))$, and $\|v^{new}(\frac{3T^{(0)}}{2})\|_{L^{2}}\leq\|v(\frac{T^{(0)}}{2})\|_{L^{2}}<c$, repeat the above steps, we get $v\in C([0,T^{*}];L^{2}(\mathbb{T}))\cap L^{2}([0,T^{*}];H^{\frac{\gamma}{2}}(\mathbb{T}))$. Since $T^{*}$ is arbitrary, we prove that $v$ is the weak global solution of (3.3), let $X$ satisfied (1.5), we prove that $v$ is the weak global solution of (1.1).
Finally, we prove that for $\gamma>\frac{5}{3}$, $v$ can be the strong global solution of (1.1). In fact, its suffice to show the uniqueness of $v$. Let $u_{0}^{1}$, $u_{0}^{2}$ belongs to $H^{\frac{1}{2}+\delta}$ for $\delta$ satisfied Proposition 3.1, and $v(t;u_{0})$ be the solution of (3.3) with $v(0)=u_{0}$. Let $v^{1}=v(t;u_{0}^{1})$, $v^{2}=v(t;u_{0}^{2})$ and $\overline{v}:=v(t;u^{1}_{0})-v(t;u^{2}_{0})$, then $\overline{v}$ satisfied the equation
$\partial_{t}\overline{v}-\Lambda^{\gamma}\overline{v}=\frac{1}{2}\partial_{x}(\overline{v}(v^{1}+v^{2}))+\partial_{x}(\overline{v}X).$
@@
$\|\overline{v}\|_{L^{2}}^{2}+\int_{0}^{t}\|\overline{v}\|_{H^{\frac{\gamma}{2}}}^{2}\mathrm{d}t^{\prime}\leq 2\int_{0}^{t}\|\overline{v}\|_{L^{2}}^{2}\mathrm{d}t^{\prime}+\int_{0}^{t}\langle\overline{v},\partial_{x}(\overline{v}(v^{1}+v^{2}))+2\partial_{x}(\overline{v}X)\rangle_{0}\mathrm{d}t^{\prime}.$
By Hölder’s inequality, Parsavel formula and Lemma 2.1, for $\gamma>\frac{5}{3}$, we have
$\langle\overline{v},\partial_{x}(\overline{v}(v^{1}+v^{2}))\rangle_{0}\leq\|\overline{v}\|_{H^{\frac{\gamma}{2}}}\|\overline{v}(v^{1}+v^{2})\|_{H^{1-\frac{\gamma}{2}}}\leq\|\overline{v}\|_{H^{\frac{\gamma}{2}}}^{2}\|(v^{1}+v^{2})\|_{H^{s}}$
for some $\frac{\gamma}{2}-1>s>1-\frac{\gamma}{2}$ which can be choose by Lemma 3.4. Similarly,
$\langle\overline{v},2\partial_{x}(\overline{v}X))\rangle_{0}\leq\|\overline{v}\|_{H^{\frac{\gamma}{2}}}^{2}\|X\|_{W^{\alpha}}.$
By Gronwall lemma, we get $\|\overline{v}\|_{L^{2}}^{2}+\int_{0}^{t}\|\overline{v}\|_{H^{\frac{\gamma}{2}}}^{2}\mathrm{d}t^{\prime}=0$, which finish our proof. ∎
@@ 3.3 Application to DP equation
This section demonstrates the application of the aforementioned theory to the study of a class of shallow water wave equations. The only difference between it and the Burgers equation lies on $\frac{3}{2}(1-\partial_{x}^{2})^{-1}\partial_{x}(u^{2})$. Note that the operator $(1-\partial_{x}^{2})^{-1}$ is an ‘stable’ operator in $H^{s}$ since
@@
Hence, we can easily get the local result for DP equation which finish the proof of Theorem 1.1 by incorporating Proposition 3.1.
For the global result, we need following energy estimate for DP equation
@@
For $\gamma\in(\frac{3}{2},2]$, let $v$ be a smooth solution of (1.12) on $[0,T]$ for $u_{0}\in H^{\frac{1}{2}+\delta}$ and $X(t)$ satisfied (1.5), then we have energy estimate as follows
$\|v\|_{L^{2}}^{2}+(2-2\nu)\int_{0}^{t}\|v\|_{H^{\frac{\gamma}{2}}}^{2}\mathrm{d}\tau\lesssim\|u_{0}\|_{L^{2}}^{2}e^{(1+C_{\nu}\|X\|_{C_{T}H^{\alpha}}^{\frac{\gamma}{\gamma-1}}+\|X\|_{C_{T}W^{\alpha}})t}+(1+C_{\nu})\|X\|_{W^{\alpha}}^{4}.$ (3.15)
@@
For the case $t\in[\delta,T]$, let $n=(1-\partial_{x}^{2})v$ and $w=(4-\partial_{x}^{2})^{-1}v$, similar as the proof of Lemma 4.1 in [CZG24], we have
$\int_{\mathbb{R}}v\partial_{x}nw+3\partial_{x}vnw\mathrm{d}x=0.$
Then we can derive to following equation
$\partial_{t}\int_{\mathbb{R}}nw\mathrm{d}x=$ $2\int_{\mathbb{R}}\langle\partial_{x}\rangle^{2}\partial_{t}vw\mathrm{d}x$
$=$ $-2\int_{\mathbb{R}}\langle\partial_{x}\rangle^{2}\Lambda^{\gamma}vw\mathrm{d}x-2\int_{\mathbb{R}}\langle\partial_{x}\rangle^{2}\partial_{x}(2vX+X^{2})w\mathrm{d}x-3\int_{\mathbb{R}}\partial_{x}(2vX+X^{2})w\mathrm{d}x.$
@@
$\int_{\mathbb{T}}\langle\partial_{x}\rangle^{2}\partial_{x}(vX)w\mathrm{d}x=$ $\int_{\mathbb{T}}\frac{1+|\xi|^{2}}{4+|\xi|^{2}}\mathcal{F}(\partial_{x}(vX))\mathcal{F}(v)\mathrm{d}x$
$\lesssim$ $\|v\|_{H^{\frac{\gamma}{2}}}\|vX\|_{H^{1-\frac{\gamma}{2}}}$
$\lesssim$ $\|v\|_{H^{\frac{\gamma}{2}}}^{\frac{2}{\gamma}}\|v\|_{L^{2}}^{2-\frac{2}{\gamma}}\|X\|_{\mathcal{C}^{\alpha}}$
$\lesssim$ $\frac{\nu}{3}\|v\|_{H^{\frac{\gamma}{2}}}^{2}+C_{\nu}\|v\|_{L^{2}}^{2}\|X\|_{W^{\alpha}}^{\frac{\gamma}{\gamma-1}},$
where we use the fact $\alpha>{1-\frac{\gamma}{2}}$. Similarly, use Lemma 2.1, we have
$\int_{\mathbb{T}}\langle\partial_{x}\rangle^{2}\partial_{x}X^{2}w\mathrm{d}x\lesssim$ $\frac{\nu}{3}\|v\|_{H^{\frac{\gamma}{2}}}^{2}+C_{\nu}\|X\|_{W^{\alpha}}^{4},$
$\int_{\mathbb{T}}\partial_{x}(2vX+X^{2})w\mathrm{d}x\lesssim$ $\frac{\nu}{3}\|v\|_{H^{\frac{\gamma}{2}}}^{2}+C_{\nu}\|v\|_{L^{2}}^{2}\|X\|_{W^{\alpha}}^{\frac{\gamma}{\gamma-1}}+C_{\nu}\|X\|_{W^{\alpha}}^{4}.$
Summarize all, by Gronwall’s inequality, we finish our proof. ∎
Similar as the proof of Lemma 3.4, we can improve the estimate in higher regularity as following, which we omit the proof.
@@
Under the assumption in Proposition 3.3, for $v\in C_{T}L^{2}\cap L^{2}_{T}H^{\frac{\gamma}{2}}$, we have energy estimate as follows
$\|\Lambda^{s}v\|_{L^{2}}+2(1-\nu)\int_{0}^{t}\|\Lambda^{s}v\|_{H^{\frac{\gamma}{2}}}^{2}\mathrm{d}t^{\prime}\leq$ $\|u_{0}\|^{2}_{L^{2}}e^{C_{\nu}\int_{0}^{t}\|v\|_{H^{\frac{\gamma}{2}}}^{2}\mathrm{d}t^{\prime}}$ (3.16)
$+\int_{0}^{t}e^{C_{\nu}\int_{t^{\prime}}^{t}\|v\|_{H^{\frac{\gamma}{2}}}^{2}\mathrm{d}t^{\prime\prime}}(\|X\|_{W^{\alpha}}^{2}\|v\|_{H^{\frac{\gamma}{2}}}^{2}+\|X\|_{W^{\alpha}}^{4})\mathrm{d}t^{\prime}$

### v2

@@ 3.2 Well-posedness
be hold for some $\sigma,s>0$, $p>1$. Observing that $\partial_{t}v=\Lambda^{\gamma}v+\partial_{x}(v^{2})+2\partial_{x}(vX)+\partial_{x}(X^{2})$. By Lemma 2.1 and Proposition 3.2, we have $\partial_{t}v\in L_{T}^{2}H^{-\sigma}$ for any $\sigma\geq 1$. The second condition of (3.13) is a consequence of Lemma (3.4). Then we prove $v$ is the solution of (3.3) which prove our claim.
To obtain the global solution, we proceed by our induction argument. Fix $\alpha\in(\frac{1}{4},\frac{1}{2})$, $\bar{T}\in[0,\infty)$. Choosing $T^{(0)}\leq\bar{T}\wedge T^{*}$ sufficiently small so that the right hand side of (3.2) is bounded by $\|u_{0}\|_{L^{2}}$($T^{(0)}$ can be only dependent on $\|X\|_{C_{\bar{T}}W^{\alpha}}$ and $\|u_{0}\|_{L^{2}}$). Consider equation (3.3) starting from $v(\frac{T^{(0)}}{2})$. Setting $v^{(1)}=v(\frac{T^{(0)}}{2}+t)$, $X^{(1)}=X(\frac{T^{(0)}}{2}+t)$, we find that $v^{(1)}(t)$ satisfies the equation
$\partial_{t}v^{(1)}(t)-\Lambda^{\gamma}v^{(1)}(t)=-\partial_{x}((v^{(1)}+X^{(1)})^{2}),\quad v^{(1)}(0)=v(\frac{T^{(0)}}{2}),$ (3.14)
then we can get a solution $v^{(1)}$ of equation on $0,T^{(0)}\wedge\bar{T}]$ by Proposition 3.1 since $\|v(\frac{T^{(0)}}{2})\|_{L^{2}}\leq\|u_{0}\|_{L^{2}}$ and for any $t\in[0,T^{*}-\frac{3T^{(0)}}{2}\wedge\bar{T}]$, we have
$\|X^{(1)}\|_{C_{t}W^{\alpha}}=\|X\|_{C_{[\frac{T^{(0)}}{2},\frac{T^{(0)}}{2}+t]}W^{\alpha}}\leq\|X\|_{C_{\bar{T}}W^{\alpha}}.$
Repeating such process, we can get $v^{(n)},X^{(n)}$ satisfy for any $t\in[0,\bar{T}-\frac{n+2}{2}T^{(0)}\wedge\bar{T}]$,
$\|v^{(n)}(\frac{T^{0}}{2})\|_{L^{2}}\leq\|u_{0}\|_{L^{2}},\quad\|X^{(n)}\|_{C_{t}W^{\alpha}}\leq\|X\|_{C_{T^{*}}W^{\alpha}},\quad v^{(n)}(t)=v^{(n-1)}(\frac{T^{0}}{2}+t),$ (3.15)
where $v^{(n)}$ on $0,T^{(0)}]$ is also a solution of (3.3). Setting $v:=v^{(0)}$ on $0,\frac{T^{(0)}}{2}]$, $v^{new}:=v^{(i)}$ on $\frac{iT^{(0)}}{2},\frac{(i+1)T^{(0)}}{2}]$ for $i\leq n$, then $v^{new}$ is a solution in $C([0,\frac{(n+1)T^{(0)}}{2}];L^{2}(\mathbb{T}))\cap L^{2}([0,\frac{(n+1)T^{(0)}}{2}];H^{\frac{\gamma}{2}}(\mathbb{T}))$, and $\|v^{new}(\frac{(n+1)T^{(0)}}{2})\|_{L^{2}}\leq\|v(\frac{T^{(0)}}{2})\|_{L^{2}}$. Choosing $n=[\frac{2\bar{T}}{T^{(0)}}]$, we get $v\in C([0,\bar{T}];L^{2}(\mathbb{T}))\cap L^{2}([0,\bar{T}];H^{\frac{\gamma}{2}}(\mathbb{T}))$. Since $\bar{T}$ is arbitrary, we obtain $v$ as a weak global solution of (3.3), since $X$ satisfy (1.5), we prove that $v$ is the weak global solution of (1.1).
Finally, we prove that for $\gamma>\frac{5}{3}$, $v$ can be the strong global solution of (1.1). In fact, it’s suffice to show the uniqueness of $v$. Let $u_{0}^{1}$, $u_{0}^{2}$ belongs to $H^{\frac{1}{2}+\delta}$ for $\delta$ satisfying Proposition 3.1, and $v(t;u_{0})$ be the solution of (3.3) with $v(0)=u_{0}$. Let $v^{1}=v(t;u_{0}^{1})$, $v^{2}=v(t;u_{0}^{2})$ and $\overline{v}:=v(t;u^{1}_{0})-v(t;u^{2}_{0})$, then $\overline{v}$ satisfies the equation
$\partial_{t}\overline{v}-\Lambda^{\gamma}\overline{v}=\partial_{x}(\overline{v}(v^{1}+v^{2}))+2\partial_{x}(\overline{v}X).$
@@
$\|\overline{v}(t)\|_{L^{2}}^{2}+\int_{0}^{t}\|\overline{v}(t^{\prime})\|_{H^{\frac{\gamma}{2}}}^{2}\mathrm{d}t^{\prime}\leq 2\int_{0}^{t}\|\overline{v}(t^{\prime})\|_{L^{2}}^{2}\mathrm{d}t^{\prime}+\int_{0}^{t}\langle\overline{v}(t^{\prime}),\partial_{x}(\overline{v}(t^{\prime})(v^{1}(t^{\prime})+v^{2}(t^{\prime})))+2\partial_{x}(\overline{v}(t^{\prime})X(t^{\prime}))\rangle_{0}\mathrm{d}t^{\prime}.$
By Hölder’s inequality, Parsavel formula and Lemma 2.1, for $\gamma>\frac{5}{3}$ and $\gamma-\frac{3}{2}>s>1-\frac{\gamma}{2}$, we have
$\langle\overline{v},\partial_{x}(\overline{v}(v^{1}+v^{2}))\rangle_{0}\leq$ $\|\overline{v}\|_{H^{\frac{\gamma}{2}}}\|\overline{v}(v^{1}+v^{2})\|_{H^{1-\frac{\gamma}{2}}}$
$\leq$ $\|\bar{v}\|^{1+\frac{1}{\gamma}+}_{H^{\frac{\gamma}{2}}}\|\bar{v}\|^{1-\frac{1}{\gamma}-}_{L^{2}}\|v^{1}+v^{2}\|_{H^{1-\frac{\gamma}{2}}}$
$\leq$ $\frac{\nu}{2}\|\bar{v}\|^{2}_{H^{\frac{\gamma}{2}}}+\|\bar{v}\|^{2}_{L^{2}}\|v^{1}+v^{2}\|^{\frac{2\gamma}{\gamma-1}+}_{H^{s}}$
which can be choose by Lemma 3.4. Similarly,
$\langle\overline{v},2\partial_{x}(\overline{v}X))\rangle_{0}\leq\frac{\nu}{2}\|\bar{v}\|^{2}_{H^{\frac{\gamma}{2}}}+\|\bar{v}\|^{2}_{L^{2}}\|X\|^{\frac{2\gamma}{\gamma-1}+}_{W^{\alpha}}.$
By Gronwall lemma, we get $\|\overline{v}(t)\|_{L^{2}}^{2}+\int_{0}^{t}\|\overline{v}(t^{\prime})\|_{H^{\frac{\gamma}{2}}}^{2}\mathrm{d}t^{\prime}\leq 0$ for any $t\geq 0$, which finish our proof. ∎
@@ 3.3 Application to DP equation
This section demonstrates the application of the aforementioned theory to the study of a class of shallow water wave equations. Taking the DP equation as an example, similar as the equation (3.3), we consider the difference equation of type of DP.
$\partial_{t}v-\Lambda^{\gamma}v=\frac{1+3(1-\partial_{x}^{2})^{-1}}{2}\partial_{x}(v^{2}+2v\tilde{X}+\tilde{X}^{2}),\quad v(0)=u_{0}.$ (3.16)
The only difference between (3.16) and (3.3) lies on $\frac{3}{2}(1-\partial_{x}^{2})^{-1}\partial_{x}(u^{2})$. Note that the operator $(1-\partial_{x}^{2})^{-1}$ is an ‘stable’ operator in $H^{s}$ since
@@
Hence, we can easily get the local result for rough DP equation. For the global result, we need the following energy estimate for DP equation.
@@
For $\gamma\in(\frac{3}{2},2]$, $\alpha\in(\frac{1}{4},\frac{\gamma}{2}-\frac{1}{2})$, $\delta$ satisfying Theorem 1.1 and $0\leq T\leq\bar{T}<\infty$. Let $v$ be a smooth solution of (3.16) on $[0,T]$ for $u_{0}\in H^{\frac{1}{2}+\delta}$ and $\tilde{X}(t)\in C_{\bar{T}}W^{\alpha}$, then we have the following energy estimate
[…]
for some small $\nu<1$.
@@
Let $n=(1-\partial_{x}^{2})v$ and $w=(4-\partial_{x}^{2})^{-1}v$, similar as the proof of Lemma 4.1 in [CZG24], by (3.16), we have
$\int_{\mathbb{T}}v\partial_{x}nw+3\partial_{x}vnw\mathrm{d}x=0.$
Then we can derive it to the following equation
$\partial_{t}\int_{\mathbb{T}}nw\mathrm{d}x=$ $2\int_{\mathbb{T}}\langle\partial_{x}\rangle^{2}\partial_{t}vw\mathrm{d}x$
$=$ $2\int_{\mathbb{T}}\langle\partial_{x}\rangle^{2}\Lambda^{\gamma}vw\mathrm{d}x+2\int_{\mathbb{T}}\langle\partial_{x}\rangle^{2}\partial_{x}(2v\tilde{X}+\tilde{X}^{2})w\mathrm{d}x+3\int_{\mathbb{T}}\partial_{x}(2v\tilde{X}+\tilde{X}^{2})w\mathrm{d}x.$
@@
$\int_{\mathbb{T}}\langle\partial_{x}\rangle^{2}\partial_{x}(v\tilde{X})w\mathrm{d}x=$ $\int_{\mathbb{T}}\frac{1+|\xi|^{2}}{4+|\xi|^{2}}\mathcal{F}(\partial_{x}(v\tilde{X}))\mathcal{F}(v)\mathrm{d}x$ (3.18)
$\leq$ $\|v\|_{H^{\frac{\gamma}{2}}}\|v\tilde{X}\|_{H^{1-\frac{\gamma}{2}}}$
$\leq$ $\|v\|_{H^{\frac{\gamma}{2}}}^{1+\frac{1}{\gamma}}\|v\|_{L^{2}}^{1-\frac{1}{\gamma}}\|\tilde{X}\|_{\mathcal{C}^{\alpha}}$
$\leq$ $\frac{\nu}{6}\|v\|_{H^{\frac{\gamma}{2}}}^{2}+C_{\nu}\|v\|_{L^{2}}^{2}\|\tilde{X}\|_{W^{\alpha}}^{\frac{2\gamma}{\gamma-1}},$
where we use the fact $\alpha>{1-\frac{\gamma}{2}}$. Similarly, by Lemma 2.1, we have
$\int_{\mathbb{T}}\langle\partial_{x}\rangle^{2}\partial_{x}\tilde{X}^{2}w\mathrm{d}x\leq$ $\frac{\nu}{6}\|v\|_{H^{\frac{\gamma}{2}}}^{2}+C_{\nu}\|\tilde{X}\|_{W^{\alpha}}^{4},$ (3.19)
$\int_{\mathbb{T}}\partial_{x}(2v\tilde{X}+\tilde{X}^{2})w\mathrm{d}x\leq$ $\frac{\nu}{3}\|v\|_{H^{\frac{\gamma}{2}}}^{2}+C_{\nu}\|v\|_{L^{2}}^{2}\|\tilde{X}\|_{W^{\alpha}}^{\frac{\gamma}{\gamma-1}}+C_{\nu}\|\tilde{X}\|_{W^{\alpha}}^{4}.$ (3.20)
Combining above estimates, we have
$\int_{\mathbb{T}}n(t)w(t)\mathrm{d}x$ $+(2-2\nu)\int_{0}^{t}\|v(t^{\prime})\|_{H^{\frac{\gamma}{2}}}^{2}\mathrm{d}t^{\prime}$ (3.21)
$\leq$ $\int_{\mathbb{T}}n(0)w(0)\mathrm{d}x+C_{\nu}\int_{0}^{t}\|v(t^{\prime})\|_{L^{2}}^{2}\|\tilde{X}(t^{\prime})\|_{W^{\alpha}}^{\frac{2\gamma}{\gamma-1}}\mathrm{d}t^{\prime}+C_{\nu}\int_{0}^{t}\|\tilde{X}(t^{\prime})\|_{W^{\alpha}}^{4}\mathrm{d}t^{\prime}.$
Noting that
$\|v(t)\|_{L^{2}}^{2}=\|\widehat{v}(t)\|_{L^{2}}^{2}\leq 4\int_{\mathbb{T}}\frac{1+\xi^{2}}{4+\xi^{2}}|\widehat{v}(t)|^{2}\mathrm{d}x=4\int_{\mathbb{T}}n(t)w(t)\mathrm{d}x\leq 4\|v(t)\|_{L^{2}}^{2}.$ (3.22)
Particularly, $\int_{\mathbb{T}}n(0)w(0)\mathrm{d}x\leq 4\|u_{0}\|_{L^{2}}^{2}$, by (3.21) we have
$\|v(t)\|_{L^{2}}^{2}+(8-8\nu)\int_{0}^{t}\|v(t^{\prime})\|_{H^{\frac{\gamma}{2}}}^{2}\mathrm{d}t^{\prime}\leq 16\|u_{0}\|_{L^{2}}^{2}+C_{\nu}\int_{0}^{t}\|v(t^{\prime})\|_{L^{2}}^{2}\|\tilde{X}(t^{\prime})\|_{W^{\alpha}}^{\frac{2\gamma}{\gamma-1}}\mathrm{d}t^{\prime}+C_{\nu}\int_{0}^{t}\|\tilde{X}(t^{\prime})\|_{W^{\alpha}}^{4}\mathrm{d}t^{\prime}.$ (3.23)
for $\nu<1$ and $t\in[0,T]$, by Gronwall’s inequality, we finish our proof. ∎
Similar as the proof of Lemma 3.4, we can improve the estimate in higher regularity as follows, which we omit the proof.
@@
Let $\gamma\in(\frac{3}{2},2]$, $\alpha\in(s+1-\frac{\gamma}{2},\frac{\gamma}{2}-\frac{1}{2})$ , $s<\gamma-\frac{3}{2}$, $\delta$ satisfy Theorem 1.1 and $0\leq T\leq\bar{T}<\infty$, let $v$ be the smooth solution of (3.16) on $[0,T]$ for $u_{0}\in H^{\frac{1}{2}+\delta}$ with $\tilde{X}\in C_{\bar{T}}W^{\alpha}$. For $t\in[0,T]$, we have the following energy estimate
$\|\Lambda^{s}v(t)\|_{L^{2}}+2(1-\nu)\int_{0}^{t}\|\Lambda^{s}v(t^{\prime})\|_{H^{\frac{\gamma}{2}}}^{2}\mathrm{d}t^{\prime}\leq$ $\|u_{0}\|^{2}_{L^{2}}e^{C_{\nu}\int_{0}^{t}\|v(t^{\prime})\|_{H^{\frac{\gamma}{2}}}^{2}\mathrm{d}t^{\prime}}+\|\tilde{X}\|_{C_{\bar{T}}W^{\alpha}}^{4}\int_{0}^{t}e^{C_{\nu}\int_{t^{\prime}}^{t}\|v(t^{\prime\prime})\|_{H^{\frac{\gamma}{2}}}^{2}\mathrm{d}t^{\prime\prime}}\mathrm{d}t^{\prime}$ (3.24)
$+\|\tilde{X}\|_{C_{\bar{T}}W^{\alpha}}^{2}\int_{0}^{t}e^{C_{\nu}\int_{t^{\prime}}^{t}\|v(t^{\prime\prime})\|_{H^{\frac{\gamma}{2}}}^{2}\mathrm{d}t^{\prime\prime}}\|v(t^{\prime})\|_{H^{\frac{\gamma}{2}}}^{2}\mathrm{d}t^{\prime},$